dukedadon:

2 over 3 x -4=12 solve.

Extrinix:

Add four to both sides \(\dfrac{2}{3}x-4=12\) \(\dfrac{2}{3}x=16\) Then, multiply by the reciprocal on both sides and you have solved for \(x\) \(\dfrac{3}{2} \times \dfrac{2}{3}x=16 \times \dfrac{3}{2}\) \(x=24\)

dukedadon:

@extrinix wrote:
Add four to both sides \(\dfrac{2}{3}x-4=12\) \(\dfrac{2}{3}x=16\) Then, multiply by the reciprocal on both sides and you have solved for \(x\) \(\dfrac{3}{2} \times \dfrac{2}{3}x=16 \times \dfrac{3}{2}\) \(x=24\)
omg, ty

Extrinix:

No problem!